A ball is dropped vertically off a 45m building. The function that describes the height of the ball as a function of time is h(t)=-5t^2+45 where height is measured in meters and time in seconds.
A) calculate the height of the ball after 2 seconds [h(2)]
B) calculate the time taken for the ball to hit the ground [h(t)=0]

Answer by addingup(3677)

A) -5t^2+45 = -5(2)^2+45 = 25 meters
-----------------------
B) The ball has to drop 45 meters (the height of the building).
Gravity: g = 9.8 m/s^2
y: vertical displacement in meters (45 meters)
t = sqrt(2gy)/g
t = sqrt(2gy)/sqrt(g^2) now simplify:
t = sqrt(2y/g)
t = sqrt 2(45)/9.8 = 3.03 seconds